OpenStudy (anonymous):

identify the oblique asymptote of f(x) = (x-1)/(4x^2+2x-3)

OpenStudy (anonymous):

y = 4x + 6 y = 4x – 2 y = 0 No Oblique Asymptote

OpenStudy (anonymous):

there will be no oblique asymptote because the degree of the denominator is greater than the degree of the numerator.

OpenStudy (anonymous):

\[f \left( x \right)=\frac{ \left( x-1 \right) }{ 4x ^{2}+2x-3 }\] like this?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

what @Data_LG2 said

OpenStudy (anonymous):

in order to get a slant or oblique asymptote, the degree of the polynomial in the numerator must be 1 more than the degree of the polynomial in the denominator. Did you give this a read? http://www.purplemath.com/modules/asymtote3.htm

OpenStudy (anonymous):

yeah i read it

OpenStudy (anonymous):

get anything from it?

OpenStudy (anonymous):

yeah, i kind of understand, its just easier for me to learn through a teacher in person, because it takees me a while to grasp things so idk

OpenStudy (anonymous):

i understand... do your best. make sure to point out which parts you're having troubles with so you can learn.

OpenStudy (anonymous):

1) degree of bottom larger that degree of the top: horizontal asymptote \(y=0\) aka the \(x\) axis 2) degrees are the same, horizontal asymptote is \(y=\text{ratio of leading coefficients}\) 3) degree of top is one more than degree of bottom: no horizontal asymptote, slant asymptote is what you get when you divide and ignore the remainder

OpenStudy (anonymous):

so is it y=0?

OpenStudy (anonymous):

because the degree on the bottom is greater than the degree on the top

OpenStudy (anonymous):

good job!

OpenStudy (ybarrap):

yes, but not oblique

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

yes, it is a horizontal asymptote.